# Code Generation with Small Language Models: A Deep Evaluation on Codeforces

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated capabilities in code generation, potentially boosting developer productivity. However, their widespread adoption remains limited by high computational costs, significant energy demands, and security risks such as data leakage and adversarial attacks. As a lighter-weight alternative, Small Language Models (SLMs) offer faster inference, lower deployment overhead, and better adaptability to domain-specific tasks, making them an attractive option for real-world applications. While prior research has benchmarked LLMs on competitive programming tasks, such evaluations often focus narrowly on metrics like Elo scores or pass rates, overlooking deeper insights into model behavior, failure patterns, and problem diversity. Furthermore, the potential of SLMs to tackle complex tasks such as competitive programming remains underexplored. In this study, we benchmark five open SLMs—Llama 3.2 3B, Gemma 2 9B, Gemma 3 12B, DeepSeek-R1 14B, and Phi-4 14B—across 280 Codeforces problems spanning Elo ratings from 800 to 2100 and covering 36 distinct topics. All models were tasked with generating Python solutions. Phi-4 14B achieved the best performance among SLMs, with a pass@3 of 63.6%, approaching the proprietary o3-mini-high (86.8%). In addition, we evaluated Phi-4 14B on C++ and found that combining outputs from both Python and C++ increases its aggregated pass@3 to 73.6%. A qualitative analysis of Phi-4 14B’s incorrect outputs revealed that some failures were due to minor implementation issues—such as handling edge cases or correcting variable initialization—rather than deeper reasoning flaws.
Index Terms:
Small Language Models, Code Generation, Competitive Programming.
@@ I Introduction
Automatic code generation has become a strategic solution to meet rising demands for developer productivity and address the ongoing shortage of skilled professionals. Recent advances in this area have progressed from early approaches like natural language-based synthesis [1] and grammar-based models [2] to large pre-trained transformers such as CodeBERT [3] and Codex [4], which have significantly advanced the state of the art.
@@
Despite these successes, benchmarks like HumanEval [4] primarily target short, simple problems [5, 6], limiting their applicability to real-world scenarios. More recent efforts have adopted competitive programming platforms such as Codeforces [7, 8], but often emphasize high-level metrics (e.g., Elo score, pass rate) without deeper analysis of problem diversity, model behavior, or failure modes. Additionally, evaluations based on offline datasets—such as those used by AlphaCode [9, 10]—may overlook key constraints like runtime and memory usage.
@@
While Large Language Models (LLMs) achieve strong performance [6, 11, 12], their high computational demands, energy usage, and security risks—such as data leakage and intellectual property exposure [13, 14, 15, 16]—can hinder adoption. In contrast, Small Language Models (SLMs) present a promising alternative: with fewer parameters, they offer lower latency, easier deployment, and greater adaptability for fine-tuning in constrained environments [17]. However, their capabilities in domains like competitive programming remain underexplored.
@@
In this study, we evaluate five general-purpose SLMs—Llama 3.2 3B, Gemma 2 9B, Gemma 3 12B, DeepSeek-R1 14B, and Phi-4 14B—on 280 Codeforces problems ranging from Elo 800 to 2100 and covering 36 topics, including Dynamic Programming, Strings, Segment Trees. All models were tasked with generating Python code, with three submissions per problem, evaluated using the Codeforces judge for correctness, runtime, and memory usage. Among the SLMs, Phi-4 14B achieved the highest performance, with a pass@3 of 63.6%. It was particularly effective on problems rated 800–1500 and in topics such as Strings (75.9%), Sorting (76.3%), and Mathematics (68.6%). It also demonstrated high semantic consistency, solving 77.5% of problems correctly within three attempts. In contrast, other SLMs—DeepSeek-R1 14B, Llama 3.2 3B, Gemma 3 12B, and Gemma 2 9B—performed significantly lower, each scoring below 24.0%.
@@
We also evaluated Phi-4 14B on C++ code generation and found that combining outputs from both Python and C++ increased its aggregated pass@3 to 73.6%. For comparison, the proprietary o3-mini-high model achieved a pass@3 of 86.8%. A manual analysis of Phi-4 14B’s incorrect outputs, supported by an expert in competitive programming, revealed that many failures were caused by minor implementation issues, or language-specific limitations, rather than fundamental reasoning flaws. These results reinforce the practical potential of small, open models as efficient and reliable alternatives for code generation, approaching the performance of state-of-the-art proprietary solutions. All data and code are publicly available online [18].
@@ II Methodology
Our study evaluates the effectiveness of open SLMs in solving programming problems commonly found on competitive programming platforms. We evaluate five open SLMs: Llama 3.2 3B [19], Gemma 2 9B [20], Gemma 3 12B [21], Phi-4 14B [22], and DeepSeek-R1 14B [23]. These models were selected based on their potential in code-centric tasks [24, 25, 23]. Additionally, they represent a diverse set of architectural designs and development teams.
To evaluate model performance, we selected 280 problems from the Codeforces, a widely used platform in education, technical interviews, and algorithmic competitions, known for its diverse and challenging problem sets. Its broad range of topics and difficulty levels makes it a strong benchmark for assessing the reasoning and code-generation capabilities of language models [26, 9, 7]. Solving Codeforces problems requires models to comprehend complex natural language descriptions, apply suitable algorithms and data structures, and produce syntactically and semantically correct implementations. Submissions are rigorously evaluated using hidden test suites that assess correctness, edge case handling, execution time, and memory usage [10]. Although the focus is on Codeforces programming problems, many of these tasks reflect real-world scenarios encountered by software engineers, involving the interpretation of requirement documents, the use of appropriate data structures, and the application of algorithms commonly used in day-to-day practice.
@@
Codeforces ranks problems using an ELO-based system, with ratings ranging from 800 to over 3000. We focused on problems rated between 800 and 2100, as fewer than 2% of active Codeforces users had ratings above 2100 at the time of writing. For each rating level within this range, we selected 20 problems based on submission popularity as of December 15, 2024. Problem statements ranged from 111 to 1,117 tokens, with an average length of 432 and a median of 404 tokens. The dataset spans 36 topics, including Implementation, Mathematics, Dynamic Programming, and Data Structures. Implementation was the most frequent tag, typically associated with problems that are conceptually simple but require intricate coding. As an example, next we show part of the “Watermelon” problem [^1] https://codeforces.com/problemset/problem/4/A illustrating the typical structure of a Codeforces task, which includes a statement, input/output format, examples, and an optional note.
Codeforces problem statements are often indirect, embedding the task within a narrative that can obscure the core requirements. This design encourages competitors to reflect, understand the problem, and identify the appropriate algorithm or data structure. Such formulations can pose additional challenges for foundation models, which must accurately interpret implicit requirements.
We developed an automated pipeline to evaluate all SLMs. Each model submitted three solutions per problem. All executions were performed locally in March 2025 using the Ollama framework [27] on a NVIDIA RTX 3060 GPU (12GB VRAM), with default parameters from the LangChain Ollama API [28] to ensure consistency across models. The process was fully automated: problem statements were extracted, prompts generated, code produced by the models, and solutions submitted to Codeforces.
To improve solution accuracy, we designed prompts with essential contextual information, including a persona definition, language-specific instructions (Python), and clear guidelines on input/output handling. Additionally, the prompt describes the structure of the Codeforces problem statement, ensuring the model interprets it correctly. Together with the actual problem, these components provide a well-rounded foundation to guide the code generation process.
{mdframed}
[ backgroundcolor=cyan!10, linewidth=1pt, linecolor=black, innerleftmargin=5pt, innerrightmargin=5pt, ] You are a highly skilled competitive programmer with 15 years of experience in the field.
@@ III Evaluation
We evaluate models using the pass@k metric [4], which estimates the probability that at least one of the top-k generated solutions is correct. As shown in Table I, Phi-4 14B (Python) outperforms all other models, achieving 48.3% at pass@1 and 63.6% at pass@3—nearly three times higher than the next best model, DeepSeek-R1 14B (15.5% at pass@1, 23.9% at pass@3). Gemma 3 12B follows with 16.9% and 19.6%, respectively, outperforming both Gemma 2 9B and Llama 3.2 3B, which stay below 11% at pass@3, suggesting that smaller models struggle with more complex programming tasks. We also analyzed the length of generated outputs. As shown in Table I, Phi-4 14B and Gemma 3 12B produce longer solutions, averaging 24.3 and 27.7 lines of code, with over 200 tokens on average. In contrast, DeepSeek-R1 14B generates more concise outputs (9.2 LOC, 82.2 tokens). Notably, Phi-4 14B also produces the longest individual solution (1,163 tokens).
TABLE I: Evaluation of Llama 3.2 3B (Lla), Gemma 2 9B (Gem2), Gemma 3 12B (Gem3), DeepSeek-R1 14B (DS), and Phi-4 14B (Phi) across pass@k, Codeforces topics, and semantic consistency.
Lla Gem2 Gem3 DS Phi
@@
pass@1 7.0% 8.4% 16.9% 15.5% 48.3%
pass@2 9.6% 9.8% 18.6% 21.1% 58.8%
pass@3 11.1% 10.4% 19.6% 23.9% 63.6%
@@
Implementation 17.6% 20.9% 34.1% 41.3% 83.5%
Sortings 7.9% 2.6% 10.5% 18.9% 76.3%
Strings 34.5% 37.9% 51.7% 48.3% 75.9%
Two Pointers 0.0% 4.3% 4.3% 12.5% 52.2%
Math 8.1% 7.0% 19.8% 20.5% 68.6%
Brute Force 14.0% 10.0% 32.0% 28.0% 68.0%
Combinatorics 0.0% 0.0% 11.1% 33.3% 66.7%
Greedy 10.7% 4.0% 12.0% 13.3% 65.3%
Binary Search 0.0% 2.6% 5.1% 12.2% 61.5%
Constructive Alg. 3.0% 6.1% 3.0% 6.1% 57.6%
@@
LOC_MEAN 19.5 13.4 27.7 9.2 24.3
LOC_MAX 81 40 99 88 106
MEAN_TOKEN 164.8 102.7 210.2 82.2 204.7
MAX_TOKEN 752 431 762 917 1,163
@@ III-A Types of Errors
When submitting a solution to a Codeforces problem, verdicts other than Accepted—such as Wrong Answer, Compilation Error, Time Limit Exceeded (TLE), Memory Limit Exceeded (MLE), and Runtime Error—indicate specific types of failures. TLE and MLE indicate that the submitted code fails to meet the time and memory constraints, respectively. Figure 1 presents the distribution of these errors across models. Wrong Answer is the most frequent error across all models, indicating incorrect outputs for one or more test cases. Runtime Error, often caused by invalid memory access or uninitialized variables, is the second most common failure, particularly affecting Llama 3.2 3B (27%) and Gemma 2 9B (15%). TLE is also notable, especially for Gemma 3 12B (39.3%) and Gemma 2 9B (10%), suggesting inefficiencies in the generated algorithms.
DeepSeek-R1 14B also had the lowest Wrong Answer rate (24.9%). Compilation Errors are especially high in DeepSeek-R1 14B (16.4%), while MLE is rare across all models. DeepSeek-R1 14B exhibited an unusual behavior, with 35.5% of its submissions labeled as Not Answered—a custom category for outputs that could not be submitted as valid code, such as explanations or partial snippets. This indicates a difficulty in generating structured, submission-ready solutions, which limits its usability in automated pipelines.
@@
Llama 3.2 3BGemma 2 9BGemma 3 12BPhi-4 14BDeepSeek-R1 14B0020202020404040406060606058.558.558.558.566.266.266.266.253.953.953.953.933.933.933.933.924.924.924.924.927272727151515153.63.63.63.66.46.46.46.47.67.67.67.655551010101039.339.339.339.39.49.49.49.40.40.40.40.42.52.52.52.50.50.50.50.51.41.41.41.40.20.20.20.216.416.416.416.40.20.20.20.2001.91.91.91.91.71.71.71.7000000000035.535.535.535.5Error Percentage (%)Wrong AnswerRuntime ErrorTime Limit ExceededCompilation ErrorMemory Limit ExceededNot Answered
Figure 1: Percentage of submissions that resulted in an error across 840 submissions (280 problems, each submitted three times).
III-B Model Effectiveness by Topic
@@
Each Codeforces problem is associated with one or more topics. Table I summarizes model performance across 10 of the 36 topics considered in our study, including Math and Greedy algorithms—each requiring distinct forms of algorithmic reasoning. Phi-4 14B outperforms all other models across every topic, achieving the highest pass@3 accuracy. Its performance is particularly strong in Implementation (83.5%), Sortings (76.3%), and Strings (75.9%), with some gaps over the second-best model. It also leads in more complex categories like Binary Search (61.5%) and Brute Force (68%), demonstrating generalization across a range of problem types. DeepSeek-R1 14B ranks second overall. Gemma 3 12B consistently improves over Gemma 2 9B, especially in Strings (51.7%) and Brute Force (32%), suggesting that moderate increases in model size or training can yield measurable gains. Llama 3.2 3B and Gemma 2 9B struggle on more demanding topics. For instance, in Combinatorics, Llama 3.2 3B scores 0%, while Phi-4 14B achieves 66.7%. A similar trend appears in Binary Search, Constructive Algorithms, and Two Pointers, where only Phi-4 14B surpasses 50% accuracy.
III-C Performance Breakdown by Difficulty
@@
As shown in Figure 2, Phi-4 14B (Python) consistently outperforms all open-small language models across nearly all difficulty levels. However, all SLMs show a performance decline as problem difficulty increases, highlighting their limitations with complex tasks and relative strength on easier ones. Each SLM exhibits different performance profiles. Llama 3.2 3B performs reasonably well up to rating 900, with only sporadic success beyond that. Gemma 2 9B shows moderate accuracy between 800–900, while Gemma 3 12B extends that range to around 1000, although with noticeable drops in accuracy after 1200. DeepSeek-R1 14B maintains relatively stable performance up to 1300. Phi-4 14B spans the full range up to 2100, but shows a decline beyond 1500, indicating challenges with higher complexity.
@@
0.80.911.11.21.31.41.51.61.71.81.922.10%20%40%60%80%100%Codeforces Rating (K)Correct answers (%)Phi-4 14B (Python)DeepSeek-R1 14BGemma 3 12BLlama 3.2 3BGemma 2 9Bo3-mini-highPhi-4 14B (C++)Phi-4 (C++ and Python)
Figure 2: Performance of models on Codeforces problems across difficulty levels (800–2100). The y-axis indicates the percentage of correct solutions out of 20 problems per level.
III-D Consistency Analysis

### v2

# Code Generation with Small Language Models: A Codeforces-Based Study

@@ Abstract
Large Language Models (LLMs) demonstrate capabilities in code generation, potentially boosting developer productivity. However, their adoption remains limited by high computational costs, among other factors. Small Language Models (SLMs) present a lightweight alternative. While LLMs have been evaluated on competitive programming tasks, prior work often emphasizes metrics like Elo or pass rates, neglecting failure analysis. The potential of SLMs in this space remains underexplored. In this study, we benchmark three open SLMs—Llama-3.2-3B, Gemma-3-12B, and Phi-4-14B—across 280 Codeforces problems spanning Elo ratings from 800 to 2100 and covering 36 distinct topics. All models were tasked with generating Python solutions. Phi-4-14B achieved the best SLM performance with a pass@3 of 63.6%, nearing o3-mini-high (86.8%). Combining Python and C++ outputs increased Phi-4-14B’s pass@6 to 73.6%. A qualitative analysis revealed some failures stemmed from minor implementation issues rather than reasoning flaws.
@@ I Introduction
Recent advances in code generation area have progressed from early approaches like natural language-based synthesis [11] and grammar-based models [12] to large pre-trained transformers such as CodeBERT [13] and Codex [14], which have significantly advanced the state of the art. Despite these successes, benchmarks like HumanEval [14] primarily target short, simple problems [15, 16], limiting their applicability to real-world scenarios.
@@
More recent efforts have adopted competitive programming platforms such as Codeforces [17, 18], but often emphasize high-level metrics (e.g., Elo score, pass rate) without deeper analysis of problem diversity, model behavior, or failure modes. Additionally, evaluations based on offline datasets—such as those used by AlphaCode [18]—may overlook key constraints like runtime and memory usage.
@@
While Large Language Models (LLMs) achieve strong performance  [16, 19, 20], their high computational demands, energy usage, and security risks—such as data leakage and intellectual property exposure [21]—can hinder adoption. In contrast, Small Language Models (SLMs) present a promising alternative: with fewer parameters, they offer lower latency, easier deployment, and greater adaptability for fine-tuning in constrained environments [22]. However, their capabilities in domains like competitive programming remain underexplored.
@@
In this study, we evaluate the performance of three small models—Llama-3.2-3B, Gemma-3-12B, and Phi-4-14B—on a code generation task over 280 Codeforces problems ranging from Elo 800 to 2100 and covering 36 topics. All models were tasked with generating Python code, with three submissions per problem, evaluated using the Codeforces judge for correctness, runtime, and memory usage. Phi-4-14B achieved the highest performance, with a pass@3 [14] of 63.6%. It was particularly effective on problems rated 800–1500 and in topics such as Strings (75.9%), Sorting (76.3%), and Mathematics (68.6%). It also demonstrated high semantic consistency [25], solving 77.5% of problems correctly within three attempts. In contrast, other SLMs, Llama-3.2-3B, Gemma-3-12B—performed significantly lower, each scoring below 19.6%.
@@
We also evaluated Phi-4-14B on C++ code generation and found that combining outputs from both Python and C++ increased its aggregated pass@6 to 73.6%. For comparison, o3-mini-high achieved a pass@3 of 86.8%. A manual analysis of Phi-4-14B’s incorrect outputs, supported by an expert in competitive programming, revealed that some failures were caused by minor implementation issues, or language-specific limitations, rather than fundamental reasoning flaws. These results reinforce the practical potential of small, open models as efficient and reliable alternatives for code generation, approaching the performance of state-of-the-art proprietary solutions. All artifacts are publicly available online. [^1] https://zenodo.org/records/17159912
@@ II Methodology
The goal of this study is to evaluate the effectiveness of three language models—Llama-3.2-3B, Gemma-3-12B, and Phi-4-14B—in solving programming problems. This evaluation is conducted from the perspective of researchers, specifically within the domain of competitive programming. To achieve this objective, we address the following research questions:
  1. RQ1
    To what extent open models as Llama-3.2-3B, Gemma-3-12B, and Phi-4-14B can answer programming assignments? To answer this question, the correct and incorrect responses provided by the Codeforces platform will be counted.
  2. RQ2
    What types of errors are most common in the responses generated by the SLMs? Investigate the most frequent failures, such as syntactic errors, incorrect logic, or failure to meet problem constraints.
  3. RQ3
    How does the performance of SLMs vary across different programming topics? Examine the SLMs’ performance on problems from various topics, such as graph algorithms, dynamic programming, sorting, or string manipulation.
Models Selection. The models were selected based on their potential in code-centric tasks [23]. Additionally, they represent a diverse set of architectural designs and development teams. In this work, we included models explicitly labeled as “small” by their developers, with fewer than 14 billion parameters—significantly smaller than proprietary state-of-the-art models from the GPT and Gemini families.
Models Setup. We used default parameters from the LangChain Ollama API (temperature 0.8 unless specified). The evaluated models were Phi-4-14B (14.7B parameters), Llama-3.2-3B (3.2B), and Gemma-3-12B (12.2B, temperature 1), all with Q4_K_M quantization. An exploratory test with Phi-4-14B showed that temperature 0.4 performed better, suggesting that default settings may not be optimal even for high-performing models.
Codeforces Platform Selection. We selected problems from the Codeforces, a widely used platform in education, technical interviews, and algorithmic competitions, known for its diverse and challenging problem sets. Its broad range of topics and difficulty levels makes it a strong benchmark for assessing the reasoning and code-generation capabilities of language models [24, 18, 17]. Solving Codeforces problems requires models to comprehend complex natural language descriptions, apply suitable algorithms and data structures, and produce syntactically and semantically correct implementations. Submissions are rigorously evaluated using hidden test suites that assess correctness, edge case handling, execution time, and memory usage [18]. Although the focus is on Codeforces programming problems, many of these tasks reflect real-world scenarios encountered by software engineers, involving the interpretation of requirement documents, the use of appropriate data structures, and the application of algorithms commonly used in day-to-day practice.
Dataset. To assess the performance of the models, we curated a dataset with 280 programming problems from the Codeforces API. Codeforces ranks problems using an ELO-based system, with ratings ranging from 800 to over 3000. We focused on problems rated between 800 and 2100, as fewer than 2% of active Codeforces users had ratings above 2100 at the time of writing. For each rating level within this range, we selected 20 problems based on submission popularity as of December 15, 2024. Problem statements ranged from 111 to 1,117 tokens, with an average length of 432.
The dataset spans 36 topics. The most frequent tags are Implementation (91 problems), Math (86), Dynamic Programming (81), Greedy (75), and Brute Force (50). Implementation problems are typically conceptually simple but require intricate coding. As an example, next we show part of the “Watermelon” problem [^2] https://codeforces.com/problemset/problem/4/A illustrating the typical structure of a Codeforces task, which includes a statement, input/output format, examples, and an optional note.
@@ 3. RQ3
Prompt Design. To improve solution accuracy, we designed prompts with essential contextual information, including a persona definition, language-specific instructions (Python), and clear guidelines on input/output handling. Additionally, the prompt describes the structure of the Codeforces problem statement, ensuring the model interprets it correctly. Together with the actual problem, these components provide a well-rounded foundation to guide the code generation process.
You are a highly skilled competitive programmer with 15 years of experience in the field.
@@
[…]
Procedure. We developed an automated pipeline to evaluate all the models. Each model submitted three solutions per problem. All executions were performed locally in March 2025 using the Ollama framework on a NVIDIA RTX 3060 GPU (12GB VRAM), with default parameters from the LangChain Ollama API to ensure consistency across models. The process was fully automated: problem statements were extracted, prompts generated, code produced by the models, and solutions submitted to Codeforces.
@@ III Evaluation
This section presents the results for each research question and provides a discussion of the findings.
III-A RQ1. Models Performance
We evaluate models using the pass@k metric [14], which estimates the probability that at least one of the top-k generated solutions is correct. As shown in Table I, Phi-4-14B (Python) outperforms all other models, achieving 48.3% at pass@1 and 63.6% at pass@3. Gemma-3-12B follows with 16.9% and 19.6%, respectively, outperforming Llama-3.2-3B, which stay below 11% at pass@3, suggesting that smaller models struggle with more complex programming tasks. We also analyzed the length of generated outputs. As shown in Table I, Phi-4-14B and Gemma-3-12B produce longer solutions, averaging 24.3 and 27.7 lines of code, with over 200 tokens on average. Notably, Phi-4-14B also produces the longest individual solution (1,163 tokens).
TABLE I: Evaluation of Llama-3.2-3B (Lla), Gemma-3-12B (Gem3), and Phi-4-14B (Phi) across pass@k, Codeforces topics, and semantic consistency.
Lla Gem3 Phi
@@
pass@1 7.0% 16.9% 48.3%
pass@2 9.6% 18.6% 58.8%
pass@3 11.1% 19.6% 63.6%
@@
Implementation 17.6% 34.1% 83.5%
Sortings 7.9% 10.5% 76.3%
Strings 34.5% 51.7% 75.9%
Two Pointers 0.0% 4.3% 52.2%
Math 8.1% 19.8% 68.6%
Brute Force 14.0% 32.0% 68.0%
Combinatorics 0.0% 11.1% 66.7%
Greedy 10.7% 12.0% 65.3%
Binary Search 0.0% 5.1% 61.5%
Constructive Alg. 3.0% 3.0% 57.6%
@@
LOC_MEAN 19.5 27.7 24.3
LOC_MAX 81 99 106
MEAN_TOKEN 164.8 210.2 204.7
MAX_TOKEN 752 762 1,163
@@
When submitting a solution to a Codeforces problem, verdicts other than Accepted—such as Wrong Answer, Compilation Error, Time Limit Exceeded (TLE), Memory Limit Exceeded (MLE), and Runtime Error—indicate specific types of failures. TLE and MLE indicate that the submitted code fails to meet the time and memory constraints, respectively. Figure 1 presents the distribution of these errors across models. Wrong Answer is the most frequent error across all models, indicating incorrect outputs for one or more test cases. Runtime Error, often caused by invalid memory access or uninitialized variables, is the second most common failure, particularly affecting Llama-3.2-3B (27%). TLE is also notable, especially for Gemma-3-12B (39.3%), suggesting inefficiencies in the generated algorithms.
@@
Llama 3.2 3BGemma 3 12BPhi-4 14B020204040606058.558.553.953.933.933.927273.63.66.46.45539.339.39.49.42.52.51.41.40.20.20.20.21.91.91.71.7Error Percentage (%)Wrong AnswerRuntime ErrorTime Limit ExceededCompilation ErrorMemory Limit Exceeded
Figure 1: Error rate across 840 submissions (280 problems × 3).
III-C RQ3. Model Effectiveness by Topic
@@
Each Codeforces problem is associated with one or more topics. Table I summarizes model performance across 10 of the 36 topics considered in our study, including Math and Greedy algorithms—each requiring distinct forms of algorithmic reasoning. Phi-4-14B outperforms all other models across every topic, achieving the highest pass@3 accuracy. Its performance is particularly strong in Implementation (83.5%), Sortings (76.3%), and Strings (75.9%), with some gaps over the second-best model. It also leads in more complex categories like Binary Search (61.5%) and Brute Force (68%), demonstrating generalization across a range of problem types. Gemma-3-12B consistently improves over Llama-3.2-3B, especially in Strings (51.7%) and Brute Force (32%), suggesting that moderate increases in model size or training can yield measurable gains. Llama-3.2-3B struggle on more demanding topics. For instance, in Combinatorics, Llama-3.2-3B scores 0%, while Phi-4-14B achieves 66.7%. A similar trend appears in Binary Search, Constructive Algorithms, and Two Pointers, where only Phi-4-14B surpasses 50% accuracy.
III-D Performance Breakdown by Difficulty
@@
Phi-4-14B (Python) consistently outperforms all open-SLMs across nearly all difficulty levels. However, all SLMs show a performance decline as problem difficulty increases, highlighting their limitations with complex tasks and relative strength on easier ones. Each SLM exhibits different performance profiles. Llama-3.2-3B performs reasonably well up to rating 900, with only sporadic success beyond that. Gemma-3-12B extends that range to around 1000, although with noticeable drops in accuracy after 1200. Phi-4-14B spans the full range up to 2100, but shows a decline beyond 1500, indicating challenges with higher complexity.
@@
0.80.911.11.21.31.41.51.61.71.81.922.10%20%40%60%80%100%Codeforces Rating (K)Correct answers (%)Phi-4-14B (Python)Gemma-3-12BLlama-3.2-3Bo3-mini-highPhi-4-14B (C++)Phi-4 (C++ and Python)
Figure 2: Performance by Codeforces difficulty (800–2100). Values show correct solutions (out of 20) with pass@3, except Phi-4-14B (C++/Python) using pass@6.
III-E Consistency Analysis